Circular No. 8449

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 X1 (LINEAR) A. Milner, Lincoln Laboratory, reports the discovery by LINEAR of

a comet with an apparent tail in p.a. 90^{o} (discovery observation below). Following posting on the ‘NEO Confirmation Page’, other observers have confirmed the object’s cometary nature from CCD images, including E. J. Christensen at Catalina (0.68-m Schmidt telescope, Dec. 9.10–9.11 UT; coma diameter ∼ 8′′ with red mag 16.2–16.6 and faint 20′′ tail in p.a. 60^{o}) and M. Tichý, M. Kočer, and J. Tichá at Kleť (1.06-m KLENOT telescope, Dec. 9.70; diffuse with coma diameter 25′′ and a wide tail in p.a. 70^{o}).

2004 UT  α_{2000}  δ_{2000}  Mag.
Dec. 7.07585  21^{h}56^{m}11.^{s}20  −4^{o}32′40′′.1  18.7

The available astrometry, preliminary parabolic orbital elements (T = 2004 Nov. 7.37 TT, q = 0.8103 AU, ω = 355^{o}.10, Ω = 0^{o}.18, i = 5^{o}.67, equinox 2000.0), and an ephemeris appear on MPEC 2004-X31. It is possible that this comet is of short period.

SUPERNOVA 2004gc IN ARP 327 M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center

for Astrophysics; and T. Matheson, National Optical Astrononmy Observatory, report that a spectrogram (range 350–740 nm) of SN 2004gc (cf. IAUC 8442), obtained by M. Calkins on Dec. 8.42 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a supernova of type Ia, with a spectral-feature age (Riess et al. 1997, A.J. 114, 722) of ∼ 3 weeks past maximum brightness. Adopting the NED recession velocity of 9620 km/s for the host galaxy, the supernova expansion velocity, derived from the minimum of Si II (rest 635.5 nm), is ∼ 9400 km/s.

V2540 OPHIUCHI T. Ak, Istanbul University; A. Retter, Pennsylvania State University;

and A. Liu, Exmouth, W. Australia, report that unfiltered CCD photometric observations were made of V2540 Oph using a 0.3-m telescope at Exmouth over 26 nights between May 2003 and June 2004. The analysis suggests a periodic signal with an amplitude of ∼ 0.02 mag. The ephemeris is T_{min} = HJD 2453151.3098 (± 0.0062) + 0.284750 (± 0.000008)E. The periodicity very likely represents the binary period.

2004 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green